# Tailoring Personality Traits in Large Language Models via Unsupervisedly-Built Personalized Lexicons

### v2

@@ 5.2 Case study
Figure 7 shows two cases demonstrating the effects of employing the UBPL method to modulate the openness and extraversion of the model. For more intriguing cases, refer to Appendix C.
Figure 7: A1: w/o UBPL; A2: w/ UBPL.
@@ 6 Conclusion
In this paper, we have introduced a novel method for tailoring the personality traits of LLMs through the utilization of custom lexicons acquired via unsupervised learning, UBPL. Unlike conventional approaches reliant on fine-tuning or prompt engineering, our method operates during the decoding phase by employing these learned custom lexicons to make subtle adjustments to the probability of the next token predicted by the original LLMs. Our method facilitates the customization of LLMs to manifest any desired combination of the Big Five personality factors in a pluggable fashion. Extensive experimentation has affirmed the effectiveness of our approach in the finer manipulation of LLMs’ personality traits. Furthermore, our method seamlessly integrates with other LLMs without necessitating updates to their parameters, demonstrating its versatility and potential for widespread application.
Acknowledgements
Appendix A the Big Five
@@ Acknowledgements
OPE-CON $+0.14$ $0.15$ $+0.20$ $0.21$ ($-0.06$, $+0.46$) $13$
OPE-EXT $+0.31$ $0.12$ $+0.43$ $0.09$ ($+0.30$, $+0.57$) $58$
OPE-AGR $+0.14$ $0.12$ $+0.21$ $0.15$ ($+0.01$, $+0.41$) $21$
OPE-NEU $-0.12$ $0.12$ $-0.17$ $0.15$ ($-0.36$, $+0.02$) $19$
CON-EXT $-0.21$ $0.15$ $+0.29$ $0.16$ ($+0.06$, $+0.52$) $21$
CON-AGR $+0.31$ $0.14$ $+0.43$ $0.12$ ($+0.26$, $+0.61$) $43$
CON-NEU $-0.32$ $0.18$ $-0.43$ $0.16$ ($-0.55$, $-0.16$) $24$
EXT-AGR $+0.18$ $0.15$ $+0.26$ $0.19$ ($+0.01$, $+0.50$) $17$
EXT-NEU $-0.26$ $0.11$ $-0.36$ $0.08$ ($-0.48$, $-0.23$) $53$
AGR-NEU $-0.26$ $0.14$ $-0.36$ $0.09$ ($-0.55$, $-0.17$) $35$
Table 2: The correlation of five personality traits. In this table, $r$ and $SD(r)$ represent the Pearson correlation coefficient and its standard deviation among the uncorrected five personality traits, $\rho$ and $SD(\rho)$ represent the corrected Spielman correlation coefficient and its standard deviation, and "Variance Due to Artifacts" describes the percentage of total variation caused by human factors in the study. (Sample size $N=144,117$ for the entire meta-analysis)
@@
Appendix B Prompt templates
The prompt templates utilized in the construction of the UBPL’s question set and answer set are depicted in Figures 8 and 9, respectively. Figure 10 illustrates the prompt template employed when assessing the degree of personality traits in the model. Furthermore, Figure 11 displays the prompt template administered to the Llama2-13b-chat model during the automatic assessment.
Figure 8: Template-1. We combined personality descriptions in "Candidate traits" into <user> prompts, and let GPT-4 generate enough SJT questions to be manually filtered to form the question set of SJTs4LLM.
Figure 9: Template-2. We combine personality descriptions in "Candidate traits" into <user> prompts, and let GPT-4 and other models generate answers containing different personality traits to form the answer set of SJTs4LLM.
Figure 10: Template-3. We use this prompt to make LLMs answer questions in SJTs4LLM.
Figure 11: Template-4. We combine "Candidate traits" and "Traits-short" into <system> and let LLMs assess the personality of an SJT question and the corresponding answer.
Appendix C More Case study
Figures 12 through 16 show specific cases of using UBPL to change the personality of LLMs. For each case, we show the SJTs question and the corresponding two answers by models (with and without UBPL), and indicate the degree of personality displayed by each answer.
Figure 12: <OPE> A1: w/o UBPL; A2: w/ UBPL.
Figure 13: <CON> A1: w/o UBPL; A2: w/ UBPL.
Figure 14: <EXT> A1: w/o UBPL; A2: w/ UBPL.
Figure 15: <AGR> A1: w/o UBPL; A2: w/ UBPL.
Figure 16: <NEU> A1: w/o UBPL; A2: w/ UBPL.
Appendix D SJTs4LLMs
To comprehensively assess the five personality traits exhibited by the subject model, a systematic approach was employed. Initially, we utilized Template-1, as detailed in Appendix B, to instruct GPT-4 in generating 400 situational judgment test (SJT) questions for each personality trait category. Following this, a meticulous manual selection process, involving de-weighting, was applied, resulting in the curation of 200 refined SJT questions for each personality trait topic. This culminated in a total of $5\times 200$ problems constituting the problem set for SJTs4LLMs.
Subsequently, Template-1 (refer to Appendix B) was employed to elicit two markedly distinct responses (High and Low) from GPT-4 and Llama2 (13b, 7b) models for each question corresponding to every personality trait topic. This process contributed to the formation of the answer set for SJTs4LLMs. The ensuing analysis delved into the content of question set subsets about the two levels of personality expression under each trait topic. To visually represent the differences between these 10 groups of answers, we use word clouds to demonstrate them, as shown in Figures 17 to 21.

### v3

@@ 5.2 Case study
Figure 7 exhibit two cases demonstrating the effects of employing the UPL method to modulate the Openness and Extraversion of Llama2-13b-chat. For more intriguing cases, refer to Appendix D.
Figure 7: A1: w/o UPL; A2: w/ UPL. The above two cases intuitively show the remarkable effect of UPL on manipulating personality expression in LLMs.
5.3 Evaluation of Model Output Quality under UPLex
To further verify that UPLex preserves the language modeling capability and fluency of large language models (LLMs), we conducted additional experiments measuring the perplexity (PPL) of model outputs across various $\beta$ values. The results are summarized in Table 2.
  Model $\backslash$ $\beta$ $(\alpha=1)$ 1.0-1.0 0.5-0.5 $0.0$ $0.5$ $1.0$
Llama2-7b-chat 6.84(0.35)6.84\,(0.35) 6.73(0.27)6.73\,(0.27) 6.81(0.24)6.81\,(0.24) 6.80(0.24)6.80\,(0.24) 6.75(0.44)6.75\,(0.44)
OpenChat3.5-7b 7.14(0.22)7.14\,(0.22) 7.20(0.34)7.20\,(0.34) 6.96(0.29)6.96\,(0.29) 7.03(0.25)7.03\,(0.25) 7.11(0.31)7.11\,(0.31)
Neural-chat-7b 13.84(0.25)13.84\,(0.25) 13.91(0.28)13.91\,(0.28) 13.70(0.28)13.70\,(0.28) 13.32(0.28)13.32\,(0.28) 13.54(0.33)13.54\,(0.33)
Baichuan2-7B-Chat 14.13(0.33)14.13\,(0.33) 15.06(0.34)15.06\,(0.34) 14.88(0.42)14.88\,(0.42) 15.01(0.52)15.01\,(0.52) 14.83(0.55)14.83\,(0.55)
Llama2-13b-chat 7.50(0.26)7.50\,(0.26) 7.10(0.36)7.10\,(0.36) 7.87(0.22)7.87\,(0.22) 7.70(0.24)7.70\,(0.24) 6.99(0.35)6.99\,(0.35)
Yi-34b-Chat 6.03(0.37)6.03\,(0.37) 6.16(0.41)6.16\,(0.41) 5.97(0.34)5.97\,(0.34) 5.85(0.35)5.85\,(0.35) 6.12(0.38)6.12\,(0.38)
Table 2: Perplexity (PPL) results of different models under various $\beta$ values $(\alpha=1)$. Values in parentheses indicate standard deviation.
As shown in Table 2, the PPL values remain stable across different $\beta$ values, indicating that UPLex has minimal impact on language model fluency and does not degrade the generative capability of the models.
@@ 6 Conclusion
In this paper, we have introduced a novel method UPLex for tailoring the personality traits of large language models (LLMs) through the utilization of custom lexicons acquired via unsupervised learning, named UPL. Unlike conventional approaches reliant on fine-tuning or prompt engineering, our method operates during the decoding phase by employing these learned custom lexicons to make subtle adjustments to the probability of the next token predicted by the original LLMs. Our method facilitates the customization of LLMs to manifest any desired combination of the Big Five personality factors in a pluggable fashion. Extensive experimentation has affirmed the effectiveness of our approach in the finer manipulation of LLMs’ personality traits. Furthermore, our method and learned lexicons can be seamlessly integrated with other LLMs without necessitating updates to their parameters, demonstrating its versatility and potential for widespread application.
Limitations
The limitations of this study are as follows:
(i) Due to the size limitation of STD we constructed, the maximum scope of UPL’s personality regulation for LLMs has not been deeply explored, which is also our future work. We also call on relevant researchers to contribute to the construction of high-quality SJT-based datasets.
(ii) We have validated our approach on models ranging in size from $7$ billion to $34$ billion, but it would be better to experiment on larger LLMs, which is what we will try to do in the future when resources allow.
Reproducibility Statement
The authors have made great efforts to ensure the reproducibility of the empirical results reported in this paper. To ensure reproducibility, we have submitted the source code of the proposed method with our paper, and plan to release the source code on GitHub upon acceptance.
Appendix A Ablation Study
Our workflow requires constructing UPL for a target LLM only once, allowing users to control the model’s personality through UPL thereafter. This high utility is one of the advantages of our method.
Additionally, when manipulating the model’s personality, the UPL is only utilized in the model’s sampling function. Since this stage merely involves dictionary mapping, the additional time cost incurred during model inference is negligible.
Apart from the quality of STD, the only potential factor influencing the effectiveness of UPL is the parameter $M$ preset during the UPL construction. This parameter, alongside the interface parameters $\alpha$ and $\beta$ provided to users, are the sole hyperparameters in our method. In this section, we discuss the impact of parameter $M$ in detail.
Figure 8: The ablation study of $M$ on Llama2-7b-chat and Baichuan2-7B-Chat. We set $\alpha$ and $\beta_{t}$ are 11 and 1-1 respectively, where tt= Extraversion. Under these settings, a lower extraversion score of the models indicates a more effective influence of UPL in modifying their personality traits. In this figure, the baseline dashed line represents the initial personality levels of the models. We observe that when $M$ is either too high or too low, the personality scores of the models after UPL adjustment tend to the baseline levels. In contrast, the lowest extraversion score is achieved when $M$ is around $0.3$. Although identifying the optimal M is challenging, it is handled by the personal-agent developers. Therefore, this does not affect the convenience and efficiency for users when utilizing UPL.
When conducting sub-word frequency analysis on STD’s answer set, we found that some subwords with strong personality representation (denoted as strong personality sub-words) tend to be ignored due to their low occurrence probability. Therefore, we execute normalization scaling for the UPL after iterative assignment. This step aims to enhance the personality scores of low-frequency strong personality sub-words while moderately reducing the personality scores of high-frequency weak personality sub-words.
As conducted in Section 3.1, we categorized the different subwords under the same personality trait into two sets based on the positive and negative values of their personality scores, namely the positive set and the negative set. Then, we transformed the personality scores of each set to make the mean personality score in each set approach our predefined parameter $M$.
Intuitively, a smaller $M$ will focus more on low-frequency strong personality sub-words, whereas a larger $M$ will focus more on high-frequency weak personality sub-words. If we do not properly balance this trade-off, it will result in suboptimal personality regulation by the UPL.
As illustrated in Figure 8, when $M$ is around $0.3$, this balance is well-maintained, resulting in a better reduction in the model’s extroversion trait.
Appendix B the Big Five
@@ Appendix A Ablation Study
OPE-CON +0.14+0.14 $0.15$ +0.20+0.20 $0.21$ (0.06-0.06, +0.46+0.46) $13$
OPE-EXT +0.31+0.31 $0.12$ +0.43+0.43 $0.09$ (+0.30+0.30, +0.57+0.57) $58$
OPE-AGR +0.14+0.14 $0.12$ +0.21+0.21 $0.15$ (+0.01+0.01, +0.41+0.41) 2121
OPE-NEU 0.12-0.12 $0.12$ 0.17-0.17 $0.15$ (0.36-0.36, +0.02+0.02) 1919
CON-EXT 0.21-0.21 $0.15$ +0.29+0.29 $0.16$ (+0.06+0.06, +0.52+0.52) 2121
CON-AGR +0.31+0.31 $0.14$ +0.43+0.43 $0.12$ (+0.26+0.26, +0.61+0.61) $43$
CON-NEU 0.32-0.32 $0.18$ 0.43-0.43 $0.16$ (0.55-0.55, 0.16-0.16) $24$
EXT-AGR +0.18+0.18 $0.15$ +0.26+0.26 $0.19$ (+0.01+0.01, +0.50+0.50) $17$
EXT-NEU 0.26-0.26 $0.11$ 0.36-0.36 $0.08$ (0.48-0.48, 0.23-0.23) $53$
AGR-NEU 0.26-0.26 $0.14$ 0.36-0.36 $0.09$ (0.55-0.55, 0.17-0.17) $35$
Table 3: The correlation of five personality traits. In this table, $r$ and $SD(r)$ represent the Pearson correlation coefficient and its standard deviation among the uncorrected five personality traits, $\rho$ and $SD(\rho)$ represent the corrected Spielman correlation coefficient and its standard deviation, and “Variance Due to Artifacts” describes the percentage of total variation caused by human factors in the study. (Sample size $N=144,117$ for the entire meta-analysis)
@@
Appendix C Prompt templates
The prompt templates utilized in the construction of the UPL’s question set and answer set are depicted in Figures 9 and 10, respectively.
Figure 11 illustrates the prompt template used in both automatic assessment and human assessment.
Furthermore, Figure 12 displays the prompt template administered to Llama2-13b-chat during the automatic assessment.
Appendix D More Case study
Figures 13 through 17 show specific cases of using UPL to change the personality of LLMs. For each case, we show the SJT question and the corresponding two answers by models (with and without UPL), and indicate the degree of personality displayed by each answer.
Appendix E STD
To comprehensively assess the five personality traits exhibited by the subject model, a systematic approach was employed. Initially, we utilized Template-1, as detailed in Appendix C, to instruct GPT-4 in generating $400$ situational judgment test (SJT) questions for each personality trait category. Following this, a meticulous manual selection process, involving de-weighting, was applied, resulting in the curation of $200$ refined SJT questions for each personality trait topic. This culminated in a total of $5\times 200$ problems constituting the problem set for STD.
Subsequently, Template-1 (refer to Appendix C) was employed to elicit two markedly distinct responses (High and Low) from GPT-4 and Llama2 (13b, 7b) models for each question corresponding to every personality trait topic. This process contributed to the formation of the answer set for STD. The ensuing analysis delved into the content of question set subsets about the two levels of personality expression under each trait topic. To visually represent the differences between these $10$ groups of answers, we use word clouds to demonstrate them, as shown in Figures 18 to 22.
Figure 9: Template-1. We combined personality descriptions in “Candidate traits” into <user> prompts, and let GPT-4 generate enough SJT questions to be manually filtered to form the question set of STD.
Figure 10: Template-2. We combine personality descriptions in “Candidate traits” into <user> prompts, and let GPT-4 and other models generate answers containing different personality traits to form the answer set of STD.
Figure 11: Template-3. We use this prompt to make LLMs answer questions in STD.
Figure 12: Template-4. We combine “Candidate traits” and “Traits-short” into <system> and let LLMs assess the personality of an SJT question and the corresponding answer.
Figure 13: <OPE> A1: w/o UPL; A2: w/ UPL.
Figure 14: <CON> A1: w/o UPL; A2: w/ UPL.
Figure 15: <EXT> A1: w/o UPL; A2: w/ UPL.
Figure 16: <AGR> A1: w/o UPL; A2: w/ UPL.
Figure 17: <NEU> A1: w/o UPL; A2: w/ UPL.